OpenStudy (emogirlatemoooocow):

A length of a rectangle is four times the width. What is the perimeter of a rectangle with width 13 inches? A. 52 inches B. 60 inches C. 65 inches D. 130 inches

OpenStudy (emogirlatemoooocow):

@tylermcmullen23

OpenStudy (welshfella):

length = 4*13 perimeter = 2( length + width)

OpenStudy (anonymous):

do 13*4 to get 52. 52+52+13+13=?

OpenStudy (emogirlatemoooocow):

130in

OpenStudy (welshfella):

right